# Turning Tabular Foundation Models into Graph Foundation Models

### v1

@@ 4.1 Datasets
The TabPFNv2 model imposes specific constraints that direct our dataset selection. In particular, TabPFNv2 is suitable for both regression tasks and multi-class classification tasks, but the latter is restricted to at most 10 classes. Additionally, it is designed for small-to-medium-scale datasets, requiring no more than roughly 10,000 training samples. As such, we select datasets that fit within these boundaries. We note that future developments in tabular foundation models may relax these constraints and enable experiments with a broader range of graph benchmarks.
@@
To comprehensively assess the capabilities of our method, we construct two collections of datasets. The first focuses on graphs with non-textual node features, which is our primary setting, while the second contains well-known graph benchmarks with text-based node features. Across these datasets, our selection covers both regression and classification tasks, and includes graphs with both homophilic and heterophilic structure, [^2] A dataset is called homophilic if edges tend to connect similar nodes (i.e., nodes with similar labels), see Newman (2003); Mironov & Prokhorenkova (2024) for details. as summarized in Table 1.
@@
While specialized graph foundation models may yield better results for text-attributed graphs, including these datasets allows us to test the generalization of our approach. Also, we explicitly exclude datasets with bag-of-words (BoW) node features, as BoW is less relevant to modern text processing, while bringing additional challenges like high-dimensionality and sparsity.
@@ 4.2 Methods
In addition to the proposed G2T-FM, we select the following baselines for evaluation.
First, traditional supervised baselines that are trained from scratch for each dataset. These include four classic GNNs: GCN (Kipf & Welling, 2017), GraphSAGE (Hamilton et al., 2017), GAT (Veličković et al., 2018), and neighborhood-attention Graph Transformer (GT) (Shi et al., 2021). [^3] This version of GT uses only local attention to the node’s neighbors, in contrast to global-attention graph transformers. For all GNNs, we use the modifications from Platonov et al. (2023) that augment GNNs with residual connections (He et al., 2016), layer normalization (Ba et al., 2016), and MLP blocks, which often significantly improve their performance. For the GraphLand datasets with tabular features, we also use LightGBM+NFA as a strong baseline — which is a popular implementation of gradient-boosted decision trees (Ke et al., 2017) with input features augmented with graph neighborhood information via NFA (Bazhenov et al., 2025). The implementation of these models closely resembles that from GraphLand (Bazhenov et al., 2025) in terms of both model architecture and hyperparameter tuning, see Appendix A for more details.
Second, we employ several publicly available graph foundation models. Despite a lot of interest in developing GFMs recently, most of the research is focused exclusively on text-attributed graphs (as discussed above), so we were able to find only a few openly available models that support node property prediction in graphs with arbitrary node feature spaces. Specifically, we employ AnyGraph (Xia & Huang, 2024), OpenGraph (Xia et al., 2024), and TS-GNN (Finkelshtein et al., 2025), which are used in an in-context learning (ICL) regime, as well as GCOPE (Zhao et al., 2024a), which is used in the finetuning (FT) regime. For all of these methods, we use the original implementations provided by the authors. Further, we were only able to evaluate these methods on datasets with node classification tasks, as none of them support node regression. [^4] While TS-GNN in theory supports node regression, its current official implementation, to the best of our knowledge, does not allow running experiments for regression tasks.
@@ 5 Experimental Results
Table 2: Evaluation results on datasets with tabular features (datasets from the GraphLand benchmark under the RL (random low) data split).
@@
LightGBM+NFA $56.34\pm 0.06$ $78.53\pm 0.01$ $46.13\pm 0.03$ $70.84\pm 0.04$ $31.71\pm 0.01$ $61.18\pm 0.03$ $60.14\pm 0.01$ $56.10\pm 0.02$ $\mathbf{3.50}$
GCN $56.27\pm 0.29$ $77.81\pm 0.14$ $44.86\pm 0.34$ $68.02\pm 0.40$ $32.00\pm 0.15$ $58.82\pm 0.24$ $75.51\pm 0.05$ $56.03\pm 0.24$ $\mathbf{4.62}$
GraphSAGE $54.43\pm 0.32$ $78.17\pm 0.09$ $45.14\pm 0.34$ $70.00\pm 0.70$ $31.44\pm 0.15$ $59.44\pm 0.26$ $66.29\pm 0.31$ $49.32\pm 0.86$ $\mathbf{5.25}$
GAT $57.41\pm 0.80$ $77.74\pm 0.20$ $45.06\pm 0.49$ $72.07\pm 1.16$ $32.63\pm 0.16$ $59.86\pm 0.19$ $72.89\pm 0.25$ $53.60\pm 0.23$ $\mathbf{3.50}$
GT $56.98\pm 0.53$ $77.34\pm 0.20$ $46.41\pm 0.68$ $69.44\pm 0.89$ $31.11\pm 0.47$ $59.55\pm 0.27$ $72.13\pm 0.13$ $53.37\pm 0.43$ $\mathbf{4.75}$
OpenGraph (ICL) $40.38\pm 1.13$ $59.09\pm 0.72$ $15.16\pm 0.83$ N/A N/A N/A N/A N/A $\mathbf{9.00}$
AnyGraph (ICL) $28.75\pm 3.56$ $63.71\pm 1.45$ $12.84\pm 0.93$ N/A N/A N/A N/A N/A $\mathbf{10.33}$
TS-GNN (ICL) $38.54\pm 0.94$ $43.46\pm 5.17$ $20.44\pm 1.05$ N/A N/A N/A N/A N/A $\mathbf{9.33}$
GCOPE (FT) $28.81\pm 1.28$ $67.16\pm 0.98$ $14.92\pm 1.56$ N/A N/A N/A N/A N/A $\mathbf{9.33}$
G2T-FM (ICL) $60.42\pm 0.27$ $77.46\pm 0.10$ $45.84\pm 0.03$ $66.68\pm 0.09$ $26.38\pm 0.07$ $60.47\pm 0.04$ $70.00\pm 0.06$ $58.75\pm 0.15$ $\mathbf{4.38}$
G2T-FM (FT) $57.65\pm 1.92$ $79.12\pm 0.21$ $47.31\pm 0.59$ $71.05\pm 0.91$ $28.52\pm 0.43$ $63.08\pm 0.28$ $74.06\pm 0.16$ $60.29\pm 0.13$ $\mathbf{2.00}$
Table 2 contains the results of the evaluation on graph datasets with tabular features, and Table 3 contains the additional results on datasets with text-derived features. In addition to results on individual datasets, we also report average ranks (AR). Below, we summarize and discuss our key observations from these results.
@@
While many GFM publications report outperforming traditional GNNs, our findings suggest otherwise. We hypothesize that this discrepancy stems from several key differences in the evaluation protocol. First, unlike some GFM studies that focus on few-shot benchmarks, we use different datasets and larger training splits (10%/10%/80%) that allow GNNs to be trained effectively from scratch. Second, we ensure our GNN baselines are highly competitive by performing a thorough hyperparameter optimization and by using the GNN architectures from Platonov et al. (2023) that include established performance-enhancing features like residual connections and normalization (Luo et al., 2024; 2025), which are often absent in simpler baselines.
@@
On datasets with tabular features, in-context performance of G2T-FM is competitive with traditional GNNs trained from scratch.
As shown in Table 2, the average rank (AR) of G2T-FM is better than that of GCN, GraphSAGE, and GT. It also achieves large gains on some datasets (e.g., tolokers-2, artnet-views). However, on other datasets (e.g., hm-prices, avazu-ctr, twitch-views) the in-context performance is weaker, indicating that finetuning is needed for the best results.
For datasets with text-based features, Table 3 shows that in-context G2T-FM is outperformed by traditional GNNs. We attribute this to TabPFNv2 being trained on tabular data: non-tabular features appear to be out-of-domain, and good performance therefore requires adaptation via finetuning.
@@
Finetuned G2T-FM outperforms traditional GNNs trained from scratch on average on both collections of datasets.
On datasets with tabular features, finetuned G2T-FM not only delivers substantial improvements on several datasets but also almost never shows significant drops relative to traditional GNNs, with the only exception being the avazu-ctr dataset.
On datasets with text-based features, finetuned G2T-FM is also strong, often bringing notable improvements over traditional GNNs. We note, however, that specialized models that process raw text more directly may achieve higher accuracy on text-attributed graphs; comparing against such models is outside the scope of this work.
Summary
These findings highlight the significant potential of graph foundation models in general, and G2T-FM in particular. To the best of our knowledge, G2T-FM is the first GFM demonstrated to be competitive with specialist GNNs trained from scratch in a setting that involves non-textual features and standard, non-few-shot dataset splits. Moreover, the superior performance of G2T-FM after finetuning indicates positive transfer from pretraining that extends beyond few-shot regimes and beyond text-attribute scenarios.
Table 3: Evaluation results on datasets with text-based features.
pubmed facebook amazon-r. questions wiki-cs AR
GCN $85.46\pm 0.18$ $91.26\pm 0.19$ $41.43\pm 0.46$ $15.42\pm 0.63$ $81.74\pm 0.20$ $\mathbf{4.00}$
GraphSAGE $86.04\pm 0.26$ $91.12\pm 0.21$ $40.07\pm 0.50$ $16.55\pm 0.61$ $81.50\pm 0.26$ $\mathbf{4.40}$
GAT $84.81\pm 0.22$ $92.61\pm 0.20$ $40.67\pm 0.53$ $16.75\pm 0.63$ $82.25\pm 0.26$ $\mathbf{3.20}$
GT $84.95\pm 0.18$ $91.71\pm 0.21$ $41.56\pm 0.38$ $14.03\pm 0.86$ $82.54\pm 0.20$ $\mathbf{3.60}$
OpenGraph (ICL) $70.30\pm 2.67$ $75.27\pm 5.05$ $29.36\pm 1.24$ $3.77\pm 0.65$ $75.66\pm 0.39$ $\mathbf{8.80}$
AnyGraph (ICL) $65.31\pm 6.26$ $61.17\pm 8.64$ $33.49\pm 3.44$ $4.27\pm 0.66$ $65.17\pm 2.51$ $\mathbf{9.00}$
TS-GNN (ICL) $64.41\pm 5.11$ $77.87\pm 2.73$ $43.00\pm 0.13$ $5.00\pm 0.48$ $46.25\pm 9.77$ $\mathbf{7.60}$
GCOPE (FT) $79.35\pm 0.70$ $85.08\pm 0.17$ $39.90\pm 0.43$ $6.59\pm 0.43$ $59.13\pm 1.20$ $\mathbf{7.60}$
G2T-FM (ICL) $88.80\pm 0.25$ $90.56\pm 0.12$ $40.63\pm 0.19$ $16.49\pm 0.16$ $76.61\pm 0.57$ $\mathbf{4.80}$
G2T-FM (FT) $90.46\pm 0.11$ $91.73\pm 0.28$ $44.71\pm 0.32$ $19.07\pm 0.53$ $79.70\pm 0.31$ $\mathbf{2.00}$
@@ 6 Ablation
Table 4: Ablation of the components of G2T-FM.“SF” stands for “Structure-based Features”, which include degree, PageRank, and Laplacian eigenvectors.
@@
First, we provide an ablation of the G2T-FM components by removing them from G2T-FM and comparing performance. Table 4 shows the results of this ablation, from which we can conclude that all the components are critical for the performance of G2T-FM. In particular, the following observation holds.
@@
Table 5: Comparison of G2T-FM against the baselines that are enhanced with the same components as G2T-FM. “(M)” stands for “Modified” and means that we add NFA, classic structure-based features and PEARL encodings to their features.
@@
G2T-FM (ICL) $60.42\pm 0.27$ $77.46\pm 0.10$ $45.84\pm 0.03$ $66.68\pm 0.09$ $26.38\pm 0.07$ $60.47\pm 0.04$ $70.00\pm 0.06$ $58.75\pm 0.15$ $\mathbf{6.62}$
G2T-FM (FT) $57.65\pm 1.92$ $79.12\pm 0.21$ $47.31\pm 0.59$ $71.05\pm 0.91$ $28.52\pm 0.43$ $63.08\pm 0.28$ $74.06\pm 0.16$ $60.29\pm 0.13$ $\mathbf{3.88}$
Second, one may argue that the performance improvements of G2T-FM come solely from the fact that it employs augmented features that are not accessible to GNN and LightGBM baselines. To verify this, we provide the baselines with exactly the same features as G2T-FM. The results are presented in Table 5.
@@
While some improvements brought by G2T-FM can be explained by its access to the features that are not used by traditional GNNs, G2T-FM shows strong performance even against the enhanced baselines. In particular, on some datasets it outperforms all other methods by a wide margin.
@@
To sum up, the gains of G2T-FM come from the synergy between the TabPFNv2 backbone and our graph-to-tabular components. The ablations show that removing any component degrades the performance, and providing the baselines with the same augmented features does not close the gap. Hence, both the backbone and the proposed components are necessary for the strong performance.

### v2

@@ 4.1 Datasets
In our experiments, we use TabPFNv2 as one of the backbones, which imposes specific constraints that direct our dataset selection. In particular, TabPFNv2 is suitable for both regression tasks and multi-class classification tasks, but the latter is restricted to at most 10 classes. Additionally, it is designed for small-to-medium-scale datasets, requiring no more than roughly 10,000 training samples. As such, we select datasets that fit within these boundaries. We note that future developments in tabular foundation models may relax these constraints and enable experiments with a broader range of graph benchmarks.
@@
To comprehensively assess the capabilities of our method, we construct two collections of datasets. The first focuses on graphs with non-textual node features (our primary setting), while the second contains well-known graph benchmarks with text-based node features. Across these datasets, our selection covers both regression and classification tasks, and includes graphs with both homophilic and heterophilic structure, [^2] A dataset is called homophilic if edges tend to connect similar nodes (i.e., nodes with similar labels), see Newman (2003); Mironov & Prokhorenkova (2024) for details. as summarized in Table 1.
@@
While specialized graph foundation models may yield better results for text-attributed graphs, including these datasets allows us to test the generalization of our approach. Also, we explicitly exclude datasets with bag-of-words (BoW) node features, as BoW is less relevant to modern text processing, while introducing additional challenges like high-dimensionality and sparsity.
@@ 4.2 Methods
In our experiments with G2T-FM, we adopt TabPFNv2 (Hollmann et al., 2025) and LimiX (Zhang et al., 2025) as backbone models. We refer to these models as G2T-TabPFNv2 and G2T-LimiX respectively. For comparison, we also evaluate the following baseline methods.
First, we include traditional supervised baselines trained from scratch for each dataset. These include four classic GNNs: GCN (Kipf & Welling, 2017), GraphSAGE (Hamilton et al., 2017), GAT (Veličković et al., 2018), and neighborhood-attention Graph Transformer (GT) (Shi et al., 2021). [^3] This version of GT uses only local attention to the node’s neighbors, in contrast to global-attention graph transformers. For all GNNs, we use the modifications from Platonov et al. (2023) that augment GNNs with residual connections (He et al., 2016), layer normalization (Ba et al., 2016), and MLP blocks, which often significantly improve their performance. For the GraphLand datasets with tabular features, we also use LightGBM+NFA as a strong baseline — which is a popular implementation of gradient-boosted decision trees (Ke et al., 2017) with input features augmented with graph neighborhood information via NFA (Bazhenov et al., 2025). The implementation of these models closely resembles that from GraphLand (Bazhenov et al., 2025) in terms of both model architecture and hyperparameter tuning, see Appendix B for more details.
Second, we employ several publicly available graph foundation models. Despite significant interest in developing GFMs recently, most of the research is focused exclusively on text-attributed graphs (as discussed above), so we were able to find only a few openly available models that support node property prediction in graphs with arbitrary node feature spaces. Specifically, we employ AnyGraph (Xia & Huang, 2024), OpenGraph (Xia et al., 2024), and TS-GNN (Finkelshtein et al., 2025), which are used in the in-context learning (ICL) regime, as well as GCOPE (Zhao et al., 2024a), which is used in the finetuning (FT) regime. For all these methods, we use the original implementations provided by the authors. Further, we were only able to evaluate these methods on datasets with node classification tasks, as none of them support node regression. [^4] While TS-GNN in theory supports node regression, its current official implementation, to the best of our knowledge, does not allow running experiments for regression tasks.
@@ 5 Experimental results
Table 2: Evaluation results on datasets with tabular features (datasets from the GraphLand benchmark under the RL (random low) data split). For each column, we highlight first, second, and third best results with a color.
@@
LightGBM+NFA $56.34\pm 0.06$ $78.53\pm 0.01$ $46.13\pm 0.03$ $70.84\pm 0.04$ $31.71\pm 0.01$ $61.18\pm 0.03$ $60.14\pm 0.01$ $56.10\pm 0.02$ $\mathbf{5.38}$
GCN $56.27\pm 0.29$ $77.81\pm 0.14$ $44.86\pm 0.34$ $68.02\pm 0.40$ $32.00\pm 0.15$ $58.82\pm 0.24$ $75.51\pm 0.05$ $56.03\pm 0.24$ $\mathbf{6.25}$
GraphSAGE $54.43\pm 0.32$ $78.17\pm 0.09$ $45.14\pm 0.34$ $70.00\pm 0.70$ $31.44\pm 0.15$ $59.44\pm 0.26$ $66.29\pm 0.31$ $49.32\pm 0.86$ $\mathbf{7.12}$
GAT $57.41\pm 0.80$ $77.74\pm 0.20$ $45.06\pm 0.49$ $72.07\pm 1.16$ $32.63\pm 0.16$ $59.86\pm 0.19$ $72.89\pm 0.25$ $53.60\pm 0.23$ $\mathbf{5.12}$
GT $56.98\pm 0.53$ $77.34\pm 0.20$ $46.41\pm 0.68$ $69.44\pm 0.89$ $31.11\pm 0.47$ $59.55\pm 0.27$ $72.13\pm 0.13$ $53.37\pm 0.43$ $\mathbf{6.62}$
OpenGraph (ICL) $40.38\pm 1.13$ $59.09\pm 0.72$ $15.16\pm 0.83$ N/A N/A N/A N/A N/A $\mathbf{11.00}$
AnyGraph (ICL) $28.75\pm 3.56$ $63.71\pm 1.45$ $12.84\pm 0.93$ N/A N/A N/A N/A N/A $\mathbf{12.33}$
TS-GNN (ICL) $38.54\pm 0.94$ $43.46\pm 5.17$ $20.44\pm 1.05$ N/A N/A N/A N/A N/A $\mathbf{11.33}$
GCOPE (FT) $28.81\pm 1.28$ $67.16\pm 0.98$ $14.92\pm 1.56$ N/A N/A N/A N/A N/A $\mathbf{11.33}$
G2T-TabPFNv2 (ICL) $60.42\pm 0.27$ $77.46\pm 0.10$ $45.84\pm 0.03$ $66.68\pm 0.09$ $26.38\pm 0.07$ $60.47\pm 0.04$ $70.00\pm 0.06$ $58.75\pm 0.15$ $\mathbf{6.38}$
G2T-TabPFNv2 (FT) $57.65\pm 1.92$ $79.12\pm 0.21$ $47.31\pm 0.59$ $71.05\pm 0.91$ $28.52\pm 0.43$ $63.08\pm 0.28$ $74.06\pm 0.16$ $60.29\pm 0.13$ $\mathbf{3.62}$
G2T-LimiX (ICL) $61.48\pm 0.30$ $77.72\pm 0.54$ $48.43\pm 0.18$ $74.96\pm 0.06$ $32.39\pm 0.14$ $64.53\pm 0.07$ $71.08\pm 0.07$ $60.95\pm 0.10$ $\mathbf{3.12}$
G2T-LimiX (FT) $61.17\pm 0.49$ $80.13\pm 0.05$ $49.88\pm 0.13$ $76.32\pm 0.17$ $33.94\pm 0.34$ $65.87\pm 0.10$ $73.16\pm 0.40$ $62.12\pm 0.10$ $\mathbf{1.38}$
Table 3: Evaluation results on datasets with text-based features. For each column, we highlight first, second, and third best results with a color.
pubmed facebook amazon-r. questions wiki-cs AR
GCN $85.46\pm 0.18$ $91.26\pm 0.19$ $41.43\pm 0.46$ $15.42\pm 0.63$ $81.74\pm 0.20$ $\mathbf{5.60}$
GraphSAGE $86.04\pm 0.26$ $91.12\pm 0.21$ $40.07\pm 0.50$ $16.55\pm 0.61$ $81.50\pm 0.26$ $\mathbf{6.00}$
GAT $84.81\pm 0.22$ $92.61\pm 0.20$ $40.67\pm 0.53$ $16.75\pm 0.63$ $82.25\pm 0.26$ $\mathbf{4.20}$
GT $84.95\pm 0.18$ $91.71\pm 0.21$ $41.56\pm 0.38$ $14.03\pm 0.86$ $82.54\pm 0.20$ $\mathbf{5.00}$
OpenGraph (ICL) $70.30\pm 2.67$ $75.27\pm 5.05$ $29.36\pm 1.24$ $3.77\pm 0.65$ $75.66\pm 0.39$ $\mathbf{10.80}$
AnyGraph (ICL) $65.31\pm 6.26$ $61.17\pm 8.64$ $33.49\pm 3.44$ $4.27\pm 0.66$ $65.17\pm 2.51$ $\mathbf{11.00}$
TS-GNN (ICL) $64.41\pm 5.11$ $77.87\pm 2.73$ $43.00\pm 0.13$ $5.00\pm 0.48$ $46.25\pm 9.77$ $\mathbf{9.60}$
GCOPE (FT) $79.35\pm 0.70$ $85.08\pm 0.17$ $39.90\pm 0.43$ $6.59\pm 0.43$ $59.13\pm 1.20$ $\mathbf{9.60}$
G2T-TabPFNv2 (ICL) $88.80\pm 0.25$ $90.56\pm 0.12$ $40.63\pm 0.19$ $16.49\pm 0.16$ $76.61\pm 0.57$ $\mathbf{6.60}$
G2T-TabPFNv2 (FT) $90.46\pm 0.11$ $91.73\pm 0.28$ $44.71\pm 0.32$ $19.07\pm 0.53$ $79.70\pm 0.31$ $\mathbf{3.00}$
G2T-LimiX (ICL) $88.96\pm 0.18$ $91.29\pm 0.14$ $44.10\pm 0.16$ $15.31\pm 0.77$ $79.99\pm 0.28$ $\mathbf{4.80}$
G2T-LimiX (FT) $89.91\pm 0.48$ $92.16\pm 0.18$ $45.67\pm 0.35$ $20.19\pm 0.30$ $82.24\pm 0.31$ $\mathbf{1.80}$
Table 2 contains the evaluation results on graph datasets with tabular features, and Table 3 contains the additional results on datasets with text-derived features. In addition to the results on individual datasets, we also report the average ranks (AR). Below, we summarize and discuss our key observations.
@@
While many GFM publications report outperforming traditional GNNs, our findings suggest otherwise. We hypothesize that this discrepancy stems from several key differences in the evaluation protocol. First, unlike some GFM studies that focus on few-shot benchmarks, we use different datasets and larger training splits (10%/10%/80%) that allow GNNs to be trained effectively from scratch. Second, we ensure our GNN baselines are highly competitive by performing a thorough hyperparameter optimization and using the GNN architectures from Platonov et al. (2023) that include established performance-enhancing features like residual connections and normalization (Luo et al., 2024; 2025), which are often absent in simpler baselines.
@@
On datasets with tabular features, G2T-FM evaluated in the in-context learning mode performs competitively with, and often better than, traditional GNNs trained from scratch.
In particular, in terms of the average rank on datasets with tabular features, G2T-LimiX outperforms all baselines trained from scratch, while G2T-TabPFNv2 performs on par with them. Results of G2T-FM on individual datasets are also strong. For example, on tolokers-2, artnet-exp, hm-prices, city-roads-M, and artnet-views, G2T-LimiX surpasses all traditional baselines, often by more than two percentage points.
At the same time, in-context performance of both G2T-TabPFNv2 and G2T-LimiX on datasets with text-based features is less impressive. In particular, G2T-TabPFNv2 is outperformed by all traditional GNNs. We attribute this to TFMs being trained on tabular data: non-tabular features appear to be out-of-domain, and good performance therefore requires adaptation via finetuning.
@@
After finetuning, G2T-FM outperforms on average all traditional baselines trained from scratch on both collections of datasets with G2T-LimiX achieving especially strong improvements.
On datasets with tabular features, G2T-LimiX achieves not only the best average rank, but also the best results on six out of eight datasets, often yielding notable improvements of more than two percentage points compared to the best GNN result.
On datasets with text-based features, finetuned G2T-FM is also strong. In particular, G2T-LimiX achieves the best average rank and often brings noticeable improvements over traditional GNNs. We note, however, that specialized models that process raw text more directly may achieve higher accuracy on text-attributed graphs. Though, comparing against such models is outside the scope of this work.
@@ 6 Ablation
In our ablation study, we focus on G2T-TabPFNv2, selecting it as the representative model because TabPFNv2 is a widely adopted and well-established tabular foundation model.
@@
Table 4: Ablation of the components of G2T-FM. SF stands for Structure-based Features, which include degree, PageRank, and Laplacian eigenvectors.
@@
First, we provide an ablation of the G2T-FM components by removing them from G2T-FM and comparing performance. Table 4 shows the results of this ablation, from which we conclude that all the components are critical for the performance of G2T-FM. In particular, the following observation holds.
@@
Table 5: Comparison of G2T-FM against the baselines that are enhanced with the same components as G2T-FM. M stands for Modified and means that we add NFA, classic structure-based features, and PEARL encodings to their features.
@@
G2T-TabPFNv2 (ICL) $60.42\pm 0.27$ $77.46\pm 0.10$ $45.84\pm 0.03$ $66.68\pm 0.09$ $26.38\pm 0.07$ $60.47\pm 0.04$ $70.00\pm 0.06$ $58.75\pm 0.15$ $\mathbf{6.62}$
G2T-TabPFNv2 (FT) $57.65\pm 1.92$ $79.12\pm 0.21$ $47.31\pm 0.59$ $71.05\pm 0.91$ $28.52\pm 0.43$ $63.08\pm 0.28$ $74.06\pm 0.16$ $60.29\pm 0.13$ $\mathbf{3.88}$
Second, one may argue that the performance improvements of G2T-FM come solely from the fact that it employs augmented features that are not accessible to the GNN and LightGBM baselines. To verify this, we provide the baselines with exactly the same features as G2T-FM. The results are presented in Table 5.
@@
While some improvements achieved by G2T-FM can be explained by its access to the features that are not used by traditional GNNs, G2T-FM shows strong performance even against the enhanced baselines. In particular, on some datasets it outperforms all other methods by a wide margin.
@@
To sum up, the gains of G2T-FM come from the synergy between the TFM backbone and our graph-to-tabular components. The ablations show that removing any component degrades the performance, and providing the baselines with the same augmented features does not close the gap. Hence, both the backbone and the proposed components are necessary for the strong performance.